# D-Wave’s Nonlinear-Program Hybrid Solver: Description and Performance Analysis

### v2

@@ V Conclusions and Further Work
Two main conclusions can be drawn from the experimentation carried out. On the one hand, the NL-Hybrid has demonstrated excellent performance compared to its competitors for those problems where the new decision variables could be used. NL-Hybrid proves to be a high-quality alternative for this type of problem, solving large instances with remarkable efficiency and obtaining solutions close to the optimum.
On the other hand, the NL-Hybrid has shown inferior performance compared to other hybrid alternatives in the MCP, where the decision variables used were of the binary type. This situation is in line with what was previously stated by D-Wave in [55]. In summary, it is a mistake to consider NL-Hybrid as a general-purpose solver, as the goal behind its implementation is not to replace the existing algorithms in HSS, but to complement them. In this way, NL-Hybrid is specifically designed to solve those problems that can take advantage of the new types of decision variables.
@@
Finally, the findings reported in this work have allowed us to identify a set of inspiring opportunities, which pave the way for several future research directions. These are some of the most interesting challenges to pursue:
@@
Study the performance of NL-Hybrid compared to other commercial hybrid methods.

### v3

@@ V Conclusions and Further Work
Given the results obtained, it is prudent to conclude that D-Wave’s NL-Hybrid emerges as a promising alternative in the realm of hybrid classical-quantum algorithms. The NL-Hybrid has demonstrated superior performance compared to its competitors, particularly in constrained problems where the new decision variables are applicable. This algorithm proves to be a high-quality approach for such complex problems, efficiently solving large instances and achieving near-optimal solutions. It is likely that NL-Hybrid will become a flagship in the field, primarily due to the innovative types of variables it incorporates, which enable the efficient handling of highly complex problems that have posed significant challenges to existing hybrid solvers.
@@
However, not everything that glitters is gold, as NL-Hybrid presents certain limitations that warrant more detailed exploration. Investigating these weaknesses will enable the scientific community to delineate the boundaries of this solver and determine the contexts in which it can perform optimally. Indeed, the NL-Hybrid has shown inferior performance compared to other hybrid alternatives in the MCP, where the decision variables used were of the binary type and the problem had no constrains. Although this is a limitation, it is not unexpected. In fact, unconstrained problems are ideally suited for the other methods, since there is no need for penalty models. This observation aligns with the statements previously made by D-Wave in [80].
Also, like any hybrid algorithm, there are further limitations that require further investigation. One of these limitations regards the overhead of asynchronous communications. Since quantum and classical computers are separate physical devices, each with its own computational interface and data transfer pipeline, the NL-Hybrid is constrained to repeatedly switch contexts and exchange intermediate data between devices. This latency prevents classical decisions from influencing the evolution of the quantum state before qubits undergo decoherence. Another aspect that requires further investigation is the scalability of the algorithm when dealing with increasingly larger problems.
With all this, the findings reported in this work have enabled us to identify a set of inspiring opportunities, paving the way for several future research directions. These are some of the most intriguing challenges to pursue:
@@
Study the performance of NL-Hybrid compared to other commercial hybrid methods, not only in terms of quality, but also in terms of scalability.